OpenStudy (anonymous):

A square has side lengths that measure 8 inches. A rectangle has a length of 10 inches and a width of 5 inches. How many more square inches are in the area of the square than the area of the rectangle?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow A(square) = 8^2 = 64 inches^2 }\) \(\Large \color{purple}{\rightarrow A(rectangle) = 10 * 5 = 50inches^2 }\) \(\Large \color{purple}{\rightarrow (64 - 50)inches^2 }\) 14 inches^2 is the difference.